Px=6a, Py=Px/a
a) Find the budget constraint equation if Bo=2(Px+Py)
b) Find the value of a if the consumer purchases 10 units of Y and 7 Units of X
1
Expert's answer
2014-08-20T12:35:19-0400
Px=6a, Py=Px/a
a)Find the budget constraint equation if Bo=2(Px+Py)
If Px = 6a, then Py = 6a/a = 6
Bo = 2(6a + 6) = 12a +12
b)Find the value of a if the consumer purchases 10 units of Y and 7 Units of X.
If X = 7, Y = 10 and the general equation ofthe budget constraint is Bo = Px*X + Py*Y, then:
12a + 12 = 6a*7 + 6*10
12a - 42a = 60 - 12
30a = -48
a = -1.6